OpenStudy (anonymous):

An adult dolphin weighs around 1410 N. How fast must he be moving as he leaves the water vertically in order to jump to a height of 1.50 m? Ignore air resistance.

OpenStudy (anonymous):

Do you know if 1.50m is the maximum height or just a height that it reached?

OpenStudy (anonymous):

well it says "to a height of 1.50 so i'm assuming that's the max height?

OpenStudy (anonymous):

http://www.physicsclassroom.com/class/1DKin/Lesson-6/Kinematic-Equations The equations are here. What you want is an equation for displacement ("jump to a height of 1.50m") List the information that you have. mg=1410N d=1.50 a=-9.80m/s^2 and if the max height is 1.50m then vf at that time=0.00m/s We have d, a, and vf. Now we find one of the equations from the link above that will allow us to determine vi. The one that I see is vf^2=vi^2+2ad. 0^2=vi^2+2(-9.8)(1.5) 29.4=vi^2 Square root both of the sides. 5.42 m/s would need to be the exit velocity necessary to propel this physics dolphin to 1.50 meters at its highest point in its flight.

OpenStudy (anonymous):

Thank you!! i totally understood how you explained it!!!

OpenStudy (anonymous):

No problem, physics has always come naturally for me. I'm glad I could help someone else understand. :)